OpenStudy (anonymous):

Can someone help me how to factor (x-4)^2=3

OpenStudy (anonymous):

@daggiemib

OpenStudy (anonymous):

take the square root of both sides \((x-4)^2=3 \\ \sqrt{(x-4)^2}=\sqrt{3}\\ (x-4)=\pm \sqrt3 \)

OpenStudy (anonymous):

okay

OpenStudy (daggiemib):

x=4+3√,4−3√

OpenStudy (anonymous):

add 4 to both sides and you get \(x=4\pm \sqrt 3 \iff 4+\sqrt3, 4-\sqrt3\)

OpenStudy (anonymous):

alright